For bromine IE/EA ratio = 10/3. Electronegativity of bromine is 2.8 on Mulliken's scale. Then it's EA in KJ/mol will be?

Grade:11

1 Answers

Khimraj
5 years ago
Given that
IE/EA = 10/3
Mulliken’s electronegativity = (IE + EA)/2 = 2.8
IE + EA = 5.6
10*EA/3 + EA = 5.6
13*EA = 16.8
so EA = 1.29 KJ/mol
Hope it clears. If you like answer then please approve it.